Can't figure out how to Calculate the Cash Flow to Stockholders (question F)

Graffiti Advertising, Inc., reported the following financial statements for the last two years. 2016 Income Statement Sales Costs of goods sold Selling and administrative Depreciation EBIT Interest EBT Taxes $566,200 274,015 124,731 54,574 $112,880 19,340 $ 93,540 37,416 56,124 $ 9,800 46,324 Net income Dividends Addition to retained earnings GRAFFITI ADVERTISING, INC Balance Sheet as of December 31, 2015 Cash Accounts receivable Inventory Current assets $ 13,340 Accounts payable 18,992 Notes payable 13,796 Current liabilities $ 9,502 14,506 $ 24,008 S 46,128 Long-term debt Net fixed assets Total assets $136,640 $ 229,966 $ 390,614 $ 344,486Owners equity 390,614 Total liabilities and owners equity GRAFFITI ADVERTISING, INC Balance Sheet as of December 31, 2016 Cash Accounts receivable Inventory Current assets $ 14,326 Accounts payable 21,097 Notes payable 22.756 Current liabilities 58,179 S 10,514 16,468 S 26,982 Long-term debt Net fixed assets Total assets $ 406,309 Owners equity $464,488 Total liabilities and owners equity S152,200 S 285,306 $464,488
21,097 Notes payable 22,756 Current liabilities 58,179 16,468 $ 26,982 Accounts receivable Inventory Current assets Long-term debt Net fixed assets Total assets $406,309Owners equity 464,488 Total liabilities and owners equity $ 152,200 $ 285,306 $ 464,488 a. Calculate the operating cash flow.(Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32 Operating cash flow 130,038 b. Calculate the change in net working capital.(Do not round intermediate calculations and round your answer to the nearest whole number Change in net working capital$ 9,079 c. Calculate the net capital spending. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32 Net capital spending $ 116,397 d. Calculate the cash flow from assets. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations ano number, e.g,32.) Cash flow from assets $ 4,562 e.Calculate the cash flow to creditors. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32 Cash flow to creditors 3,780 f. Calculate the cash flow to stockholders. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations a number, e.g., 32.) Cash flow to stockholders
Answer #1

At December 31, 2016 equity was 285,306

At the start of the year, or on Dec 31, 2015 equity was 229,966

Change in equity was = 285306 - 229966 = 55340

Also as per income statement addition to retained earnings was 46,324

Net change in equity will be = change in owner's equity - retained earnings = 55340-46324 = 9016

Cash flow to shareholders = Dividends paid - Net change in equity

=9800 - 9016

=784